# Diagnosing the AGN population origin of TeV neutrinos with their spatial correlation

### v1

@@ Abstract
The recent IceCube detection of TeV neutrinos from some nearby Seyfert galaxy (e.g., NGC 1068) suggests that active galactic nuclei (AGN) could make a significant contribution to the diffuse flux of astrophysical neutrinos. The absence of TeV gamma-rays from NGC 1068 indicates neutrino production in compact gamma-ray-opaque region. The vicinity of the supermassive black hole, such as disk-corona, is an ideal region, where the high radiation density leads to efficient neutrino production as well as the gamma-ray attenuation. Disk-corona models predict that the neutrino emission from AGNs correlates with X-ray emission, which traces the coronal activity. In this paper, we assess whether the X-ray AGN population origin for TeV neutrinos can be tested by using the spatial correlation between the neutrino population and X-ray AGN population with future neutrino telescopes. By performing simulations, we find that, the AGN origin of the neutrino background above 100 TeV can be tested at a confidence level of $\sim 2.4\sigma$ with five-year observations of IceCube-Gen2, which has an angular resolution of $\sim 0.2$ degree. With better angular resolution and sensitivity in the energy range of above $300$ TeV, a 30-$\rm km^{3}$} underwater neutrino telescope, such as High-energy Underwater Neutrino Telescope (HUNT), is expected to reach a significance of $\sim 8.6\sigma$ in testing the association after five years of exposure.
@@ I Introduction
The neutrino flux measured from NGC 1068 is more than an order of magnitude larger than the upper limits of TeV gamma-ray emission placed by MAGIC and HAWC (MAGIC1068-2019; HAWC1068-2022). The difference between the observed neutrino flux and gamma-ray flux from NGC 1068 suggests that the neutrinos must be produced in a region that is opaque to GeV–TeV gamma-rays that would otherwise accompany the neutrinos. In Seyfert galaxies like NGC 1068, a hot and magnetized corona above the disk is formed (see e.g. MillerStone2000accretion). The dense environments near the supermassive black holes together with the acceleration of cosmic rays (CRs) in the corona offer suitable conditions for the production of high energy neutrinos. Such a scenario, commonly referred to as disk-corona models, has been shown to be able to accommodate the high level of neutrino emission from the X-ray bright AGN NGC 1068 (Murase2020; Inoue2020; Kheirandish2021; Eichmann2022).
@@
The large flux contrast between neutrinos and gamma rays in NGC 1068 is in agreement with the picture of gamma-ray “hidden” sources for the origin of the isotropic neutrino background (Senno2015; Murase2016PhRvL; Bechtol2017ApJ; Capanema2020; Fang2022), as required by comparison of isotropic diffusive neutrino flux measured by IceCube (Aartsen2020PhRvL.125l1104A; ICatmneutrino2022) and the isotropic gamma-ray background observed by Fermi-LAT (Ackermann2015ApJDiffuse). This consistency further supports the main contributors of the neutrino background are “hidden” sources. Building on this framework, Fiorillo2025 extended the NGC 1068 result to a population study of X-ray–emitting AGNs and constructed an AGN luminosity sequence, demonstrating that the cumulative neutrino emission from such sources can account for the diffuse neutrino background observed by IceCube in the 11001–100 TeV energy range.
@@
However, the limited angular resolution and statistics of current IceCube data make it difficult to establish a significant spatial correlation between neutrinos and the population of X-ray AGNs. Previous searches by the IceCube Collaboration have mainly relied on stacking analyses that combine the likelihood contributions from multiple candidate sources (e.g. 2021PhRvDZhouBeiradioAGN; ic2023ApJAGNsblazars; ic2025ApJHXA), but no significant excess has yet been found so far. Future neutrino telescopes, such as IceCube-Gen2 (IceCube-gen22021), KM3NeT (KM3NET2016JPhG), Huge Underwater high-energy Neutrino Telescope (HUNT) (HUNTproposal2023), Tropical Deep-sea Neutrino Telescope (TRIDENT) (TRIDENT2023NA), NEON (NEON2025) and Baikal-GVD (Baikal2022), will have larger effective volumes and improved angular precision, and are therefore crucial for testing this connection. Motivated by this, in this work we present a cross-correlation analysis to test the X-ray AGN population origin of TeV-PeV neutrinos with future neutrino telescopes. We will use IceCube-Gen2 and HUNT as two examples for the analyses. We assume the neutrino luminosity to be positively related to the X-ray luminosity of the AGN, given that the X-ray luminosity may reflect the energy production rate in the corona, and is also proportional to the target density for photopion production. By combining the AGN X-ray luminosity distribution with the expected performance of future neutrino telescopes, we forecast the prospects for detecting such a AGN-neutrino correlation.
II Simulating Neutrino-AGN Associations: Models and Methodology
@@
In both models, we account for contamination from atmospheric neutrinos, which form an irreducible background in high-energy neutrino observations. According to the IceCube 9.5-year dataset (ICatmneutrino2022), the astrophysical signal starts to dominate the atmospheric neutrinos from 100 TeV (ICatmneutrino2022).
@@
The AGN catalog from the eROSITA Final Equatorial-Depth Survey (eFEDS) (eFEDSAGN2022) forms the basis of our analysis. It includes 22,079 AGNs selected by X-rays in a contiguous 142 deg2 field, with fluxes measured in the 0.22.30.2-2.3 keV band. Given its depth and well-characterized selection, the eFEDS sample provides a representative snapshot of the AGN population accessible to current wide-field X-ray surveys. Under the AGN-association hypothesis, we normalize the simulation by requiring that the total number of modeling neutrinos equals the expected number derived from the measured diffuse astrophysical neutrino flux plus the atmospheric neutrino background. These fluxes, taken from muon-neutrino flux with 9.5 years of IceCube data (ICatmneutrino2022), are integrated over the chosen energy threshold and then folded with the effective area of the detector to yield the expected total count of events $N_{\nu}^{\rm tot}$. We assume that a fixed fraction $f$ of these events is assigned to originate from AGNs, while the remainder is drawn from an isotropic atmospheric background. The parameter $f$ denotes the fraction of astrophysical neutrinos with energies above a given threshold relative to the total neutrino population. The probability of each AGN producing a neutrino is taken to be proportional to its X-ray flux ($F_{X}$), i.e., $P_{i}\propto F_{X,i}$, ensuring that the sum of all AGNs reproduces $f\cdot N_{\nu}^{\rm tot}$. For each AGN-origin neutrino assigned, its arrival direction is offset from the source position according to a two-dimensional Gaussian with a width equal to the detector’s point-spread function (PSF). In contrast, under the null hypothesis, neutrino positions are randomly drawn from a uniform distribution across the same sky footprint, with no relation with AGN locations.
III Unbinned Likelihood Framework for Spatial Correlation Analysis
III.1 Likelihood Method
@@
An unbinned likelihood ratio method can be used to distinguish astrophysical neutrinos from the atmospheric background produced by cosmic ray interactions (BraunML2008; crequeSeekingNeu2022; IC2024MOJAVEAGNApJ). The optimal estimator to determine the fraction $f$ of neutrinos associated with AGN is obtained using the unbinned maximum likelihood (ML) approach, formulated as
@@
$S_{i}=\sum_{\alpha}w_{\alpha}\frac{1}{2\pi\sigma_{\rm psf}^{2}}\exp\left(-\frac{(\vec{x}_{i}-\vec{x}_{\alpha})^{2}}{2\sigma_{\rm psf}^{2}}\right),$ (2)
where $\vec{x}_{i}$ and $\vec{x}_{\alpha}$ are the reconstructed positions of the neutrino and the $\alpha$-th AGN, respectively, and $w_{\alpha}=F_{\alpha}/\sum_{k}F_{k}$ reflects the X-ray flux-weighted contribution of each AGN. Here $F_{\alpha}$ denotes the observed X-ray flux of the $\alpha$-th AGN in the 0.2–2.3 keV band, as reported in the eFEDS catalog. The parameter $\sigma$ characterizes the angular resolution (i.e. PSF) of the neutrino detector. The background PDF $B_{i}$ accounts for neutrinos arising from the isotropic atmospheric background and is modeled as a uniform distribution over the analysis region: $B_{i}=\frac{1}{\Omega}$, where $\Omega$ is the solid angle subtended by the sky region covered by the AGN catalog and neutrinos.
@@
which measures the relative preference for a nonzero $f$ over the null hypothesis. To assess the uncertainty and confidence level on $f$, we define the signal-to-noise ratio as
$\frac{S}{N}_{f}\equiv\frac{\hat{f}}{\sigma_{f}},$ (4)
where $\hat{f}$ is the maximum-likelihood estimate and $\sigma_{f}$ is the standard deviation characterizing the statistical uncertainty in $f$. Instead of relying on asymptotic properties of the likelihood (Wilks’ theorem wilks1938), we compute $\sigma_{f}$ empirically by repeating the analysis on many simulated datasets with the same configuration. This Monte Carlo procedure provides a more robust estimate of the confidence interval for $f$, particularly in regimes where the number of signal events is small.
III.2 Mitigating Source Confusion in ML
Figure 1: Deviation of the recovered signal fraction from the true value, $f_{\rm recover}-0.5$, as a function of PSF and average source separation. Red (blue) regions indicate positive (negative) deviations, while the white region corresponds to no bias (i.e., $f_{\rm recover}=0.5$). The green dashed line marks $d=3\sigma_{\rm psf}$, above which the recovered fraction converges to the true value $f_{\rm recover}=0.5$.
Likelihood-based methods provide a powerful statistical framework for estimating the neutrino signal fraction $f$. However, it does not sufficiently address the impact of source confusion—a fundamental limitation in astronomical surveys that arises when the instrument’s angular resolution is too coarse to distinguish nearby sources (barcons1992). As also noted by Kowalski2025, stacking becomes unreliable when the average number of sources within a single PSF exceeds one, resulting in systematic underestimate of $f$.
@@ III.2 Mitigating Source Confusion in ML
Figure 1 summarizes the results for a fixed true value $f=0.5$. When the average separation satisfies $\bar{d}\gtrsim 3\sigma_{\rm psf}$, the maximum-likelihood estimator reliably recovers the true $f$. In contrast, for denser source populations ($\bar{d}\lesssim 3\sigma_{\rm psf}$), confusion effects cause $f$ to be systematically underestimated. This yields a practical condition for method applicability: the source population must be sufficiently sparse, such that the average spacing between sources exceeds approximately three times the PSF width. Importantly, we verify that this trend is valid for a wide range of values of $f$, not just for the specific case of $f=0.5$.
Based on this empirical criterion, we set a threshold of X-ray flux $F_{\rm min}$ for AGNs chosen in all subsequent analyses, so that chosen subset of AGNs with $F_{\rm X}\geq F_{\rm min}$ satisfy $d>3\sigma_{\rm psf}$. This ensures that the selected AGN population remains well-resolved, thereby mitigating confusion-induced bias and preserving the validity of the maximum-likelihood inference. The chosen subset therefore represents only the relatively bright portion of the whole AGN population. To infer the total neutrino contribution from the whole AGN population, we compute a correction factor based on the cumulative flux ratio between the flux-limited subset and the full sample:
$f_{\rm total}=f_{\rm bright}\cdot\frac{\sum_{\rm all}F_{X}}{\sum_{F_{X}>F_{\rm min}}F_{X}}\,$, (5)
where $f_{\rm bright}$ is the fraction inferred from the likelihood analysis of the bright subset. This scaling accounts for the contribution of unresolved faint sources that are excluded by the $F_{\rm min}$ cut. All reported values of $f$ in this work, including those shown in Figure 1, incorporate this flux-based correction to ensure consistency with the full AGN population, even though the likelihood analysis itself is restricted to the bright, confusion-free subset.
IV Predictive Power for Future Neutrino Detectors
Next-generation neutrino observatories are expected to greatly enhance the capability to detect astrophysical neutrino sources. Proposed facilities such as IceCube-Gen2, HUNT, TRIDENT, NEON and Baikal-GVD, also promise enhanced exposure and angular resolution. These advancements motivate us to examine how the improvements in angular resolution, event statistics, and energy threshold affect the prospects for diagnosing the neutrino–AGN correlation.

### v3

# Diagnosing the AGN origin of diffuse TeV neutrinos with their spatial correlation

@@ Abstract
The recent detection of TeV neutrinos from some nearby Seyfert galaxy (e.g., NGC 1068) by IceCube suggests that active galactic nuclei (AGNs) could make a significant contribution to diffuse astrophysical neutrinos. The absence of TeV gamma-rays from NGC 1068 indicates neutrino production in a compact opaque region of gamma-rays. The vicinity of the supermassive black hole, such as the disk-corona, is an ideal region, where the high radiation density leads to efficient neutrino production and gamma-ray attenuation. Disk-corona models predict that the neutrino emission from AGNs correlates with X-ray emission, which traces the coronal activity. In this paper, we assess whether diffuse TeV neutrinos can originate from X-ray–emitting AGNs with Monte Carlo simulations, considering the predicted performance of future neutrino telescopes. We test this hypothesis by searching for spatial correlations between the X-ray AGN population and high-energy neutrinos, assuming that the neutrino flux scales with the AGN X-ray flux. Based on detailed Monte Carlo simulations, we find that an AGN origin of diffuse neutrinos can be tested at a significance of about $3\sigma$ with five years of IceCube-Gen2 observations. With improved angular resolution and sensitivity, a 30 km3-scale underwater neutrino telescope (such as the High-energy Underwater Neutrino Telescope) is expected to reach a significance of about $8\sigma$ with one year of exposure. The detection significance decreases if AGNs contribute partially to the total astrophysical neutrino flux. Our results highlight the critical role of angular resolution in diagnosing the AGN origin of diffuse TeV neutrinos.
Neutrino astronomy (1100) — Active galactic nuclei (16)
@@ I Introduction
The neutrino flux measured from NGC 1068 is more than an order of magnitude larger than the upper limits of TeV gamma-ray emission placed by MAGIC and HAWC (Acciari et al., 2019; Willox and HAWC Collaboration, 2022). The difference between the observed neutrino flux and gamma-ray flux from NGC 1068 suggests that the neutrinos must be produced in a region that is opaque to GeV–TeV gamma-rays that would otherwise accompany the neutrinos. In Seyfert galaxies like NGC 1068, a hot and magnetized corona above the disk is formed (see e.g. Miller and Stone 2000). The dense environments near the supermassive black holes together with the acceleration of cosmic rays (CRs) in the corona offer suitable conditions for the production of high energy neutrinos. Such a scenario, commonly referred to as disk-corona models, has been shown to be able to accommodate the high level of neutrino emission from the X-ray bright AGN NGC 1068 (Murase et al., 2020; Inoue et al., 2020; Kheirandish et al., 2021; Eichmann et al., 2022).
@@
The large flux contrast between neutrinos and gamma rays in NGC 1068 is in agreement with the picture of gamma-ray “hidden” sources for the origin of the isotropic neutrino background (Senno et al., 2015; Murase et al., 2016; Bechtol et al., 2017; Capanema et al., 2020; Fang et al., 2022), as required by comparison of isotropic diffusive neutrino flux measured by IceCube (Aartsen et al., 2020b; Abbasi et al., 2022b) and the isotropic gamma-ray background observed by Fermi-LAT (Ackermann et al., 2015). This consistency further supports the main contributors of the neutrino background are “hidden” sources. Building on this framework, Fiorillo et al. (2025) extended the NGC 1068 result to a population study of X-ray–emitting AGNs using the luminosity function of AGNs, showing that the cumulative neutrino emission from such sources can account for the diffuse neutrinos observed by IceCube in the 1-100 TeV energy range.
@@
However, the limited angular resolution and statistics of current IceCube data make it difficult to establish a significant spatial correlation between neutrinos and the population of X-ray AGNs. Previous searches for spatial correlation have mainly relied on some classes of AGNs, such as blazars and radio-bright AGNs (e.g. Zhou et al. 2021; Abbasi et al. 2023, 2025b), but no significant excess has been found so far. Future neutrino telescopes, such as IceCube-Gen2 (Aartsen et al., 2021), KM3NeT (Adrián-Martínez et al., 2016), Huge Underwater high-energy Neutrino Telescope (HUNT) (Huang et al., 2024), Tropical Deep-sea Neutrino Telescope (TRIDENT) (Ye et al., 2023), NEON (Zhang et al., 2025) and Baikal-GVD (Belolaptikov et al., 2022), will have larger effective volumes and improved angular precision, and are therefore crucial for testing this correlation. Motivated by this, in this work, we assess whether the X-ray AGN origin for diffuse TeV neutrinos can be tested by using the spatial correlation between the X-ray AGN population and neutrino population with future neutrino telescopes. We will use IceCube-Gen2 and HUNT as two examples of future neutrino telescopes for the analyses. We assume the neutrino luminosity to be positively related to the X-ray luminosity of the AGN, given that the X-ray luminosity may reflect the energy production rate in the corona, and is also proportional to the target density for photopion production. By combining the AGN X-ray luminosity distribution with the expected performance of future neutrino telescopes, we forecast the prospects for detecting such a AGN-neutrino correlation.
II Analysis technique
II.1 Simulating Neutrino-AGN Associations
@@
In both models, we account for contamination from atmospheric neutrinos, which form an irreducible background in high-energy neutrino observations. According to the IceCube 9.5-year dataset (Abbasi et al., 2022b), the astrophysical signal starts to dominate the atmospheric neutrinos from 100 TeV. This motivates a further decomposition of the observed neutrino sample into its astrophysical and atmospheric components. We introduce the parameter $\beta$, defined as the fraction of astrophysical neutrinos relative to the total neutrino population (including atmospheric background) above a given energy threshold:
$\beta(E_{\rm th})\equiv\frac{N_{\nu}^{\rm astro}(E>E_{\nu}^{\rm th})}{N_{\nu}^{\rm all}(E>E_{\nu}^{\rm th})}.$ (1)
The value of $\beta$ depends sensitively on the chosen neutrino energy threshold $E_{\nu}^{\rm th}$ and increases rapidly toward unity at higher energies, reflecting the diminishing contribution from atmospheric neutrinos. In this work, $\beta(E_{\nu}^{\rm th})$ is derived directly from IceCube observations (Abbasi et al., 2022b) and is illustrated in Figure 1. In addition, we define the $\alpha$ as the fraction of astrophysical neutrinos produced by AGNs:
$\alpha\equiv\frac{N_{\nu}^{\rm AGN}}{N_{\nu}^{\rm astro}},$ (2)
which characterizes the efficiency of AGNs as neutrino sources relative to the total astrophysical neutrino flux.
Figure 1: The fraction of astrophysical neutrinos $\beta$ as a function of the neutrino energy threshold $E_{\nu}^{\rm th}$, derived from IceCube observations (Abbasi et al., 2022b). As the energy threshold increases, the atmospheric neutrino background is progressively suppressed, causing $\beta$ to approach unity at high energies.
The AGN catalog from the eROSITA Final Equatorial-Depth Survey (eFEDS) (Liu et al., 2022) forms the basis of our analysis. It includes 22,079 AGNs selected by X-rays in a contiguous 142 deg2 field, with fluxes measured in the 0.22.30.2-2.3 keV band. Given its depth ($6.5\times 10^{-15}{\rm erg\,s^{-1}\,cm^{-2}}$, Brunner et al. 2022) and well-characterized selection, the eFEDS sample provides a representative snapshot of the AGN population accessible to current wide-field X-ray surveys. Although eFEDS covers only a limited sky area, forthcoming all-sky and wide-area eROSITA AGN catalogs (e.g., Salvato et al., 2025) with comparable depth will substantially increase the AGN statistics and further enhance the sensitivity of our method.
@@
Under the AGN-association hypothesis, we normalize the simulation by requiring that the total number of modeling neutrinos equals the expected number derived from the measured diffuse astrophysical neutrino flux plus the atmospheric neutrino background. These fluxes, taken from muon-neutrino flux with 9.5 years of IceCube data (Abbasi et al., 2022b), are integrated over the chosen energy threshold and then folded with the effective area of the detector to yield the expected total count of events $N_{\nu}^{\rm tot}$. The probability of each AGN producing a neutrino is taken to be proportional to its X-ray flux ($F_{X}$), i.e., $P_{i}\propto F_{X,i}$. The sum of all AGN population reproduces $f_{\rm tot}\cdot N_{\nu}^{\rm tot}$.in which the fraction $f$ denotes the fraction of astrophysical neutrinos with energies above a given threshold relative to the total neutrino population. Because the AGN catalog is flux-limited, the parameter $f$ captures only the contribution from X-ray–bright AGNs. To estimate the total AGN population contribution, including X-ray–faint AGNs below the threshold $F_{\rm min}$, we extrapolate using the AGN X-ray luminosity function. The total AGN neutrino fraction is then given by
$f_{\rm tot}=f\cdot\frac{\sum_{\rm all}F_{X}}{\sum_{F_{X}>F_{\rm min}}F_{X}},$ (3)
and the expected value of this quantity is $\alpha\times\beta$.
For each AGN-origin neutrino assigned, its arrival direction is offset from the source position according to a two-dimensional Gaussian with a width equal to the detector’s point-spread function (PSF). In contrast, under the null hypothesis, neutrino positions are randomly drawn from a uniform distribution across the same sky footprint, with no relation with AGN locations.
II.2 Unbinned Likelihood Framework for Spatial Correlation Analysis
@@
An unbinned likelihood ratio method can be used to distinguish astrophysical neutrinos from the atmospheric background produced by cosmic ray interactions (Braun et al., 2008; Creque-Sarbinowski et al., 2022; Abbasi et al., 2024). The analysis is designed to measure the fractional contribution of AGN-generated neutrinos to the all-sky neutrino sample. The optimal estimator to determine the fraction $f$ of neutrinos associated with AGN is obtained using the unbinned maximum likelihood (ML) approach, formulated as
@@
$S_{i}=\sum_{j}w_{j}\frac{1}{2\pi\sigma_{\rm psf}^{2}}\exp\left(-\frac{(\vec{x}_{i}-\vec{x}_{j})^{2}}{2\sigma_{\rm psf}^{2}}\right),$ (5)
where $\vec{x}_{i}$ and $\vec{x}_{j}$ are the reconstructed positions of the neutrino and the $j$-th AGN, respectively, and $w_{j}=F_{j}/\sum_{k}F_{k}$ reflects the X-ray flux-weighted contribution of each AGN. Here $F_{j}$ denotes the observed X-ray flux of the $j$-th AGN in the 0.2–2.3 keV band, as reported in the eFEDS catalog. The parameter $\sigma_{\rm psf}$ characterizes the angular resolution (i.e. PSF) of the neutrino detector. The background PDF $B_{i}$ accounts for neutrinos arising from the isotropic atmospheric background and is modeled as a uniform distribution over the analysis region: $B_{i}=\frac{1}{\Omega}$, where $\Omega$ is the solid angle subtended by the sky region covered by the AGN catalog and neutrinos.
@@
which measures the relative preference for a nonzero $f$ over the null hypothesis. In the large-sample limit, TS approximately follows a $\chi^{2}$ distribution with one degree of freedom, and the statistical significance can be estimated as $\sigma\simeq\sqrt{\mathrm{TS}}$. The validity of this approximation has been verified with $5\times 10^{6}$ background-only Monte Carlo simulations (see Appendix A). Finally, we note that the total AGN neutrino fraction inferred from the likelihood analysis, $f_{\rm tot}$, admits a simple physical interpretation.
To quantify the uncertainty on the total AGN neutrino fraction, we define the relative error $\delta$:
$\delta\equiv\delta_{f_{\rm tot}}=\frac{\sigma_{f_{\rm tot}}}{f_{\rm tot}},$ (7)
where $\sigma_{f_{\rm tot}}$ is obtained from the distribution of $f_{\rm tot}$ over $800$ independent Monte Carlo realizations, and therefore includes both statistical fluctuations (i.e., the Poisson fluctuations in the expected number of neutrino events) and systematic uncertainties, which will be discussed in Section IV.
II.3 Mitigating Source Confusion in Analyses
Figure 2: Deviation of the recovered signal fraction from the true value, $f_{\rm tot}-0.5$, as a function of PSF and average source separation. Red (blue) regions indicate positive (negative) deviations, while the white region corresponds to no bias (i.e., $f_{\rm tot}=0.5$). The green dashed line marks $d=3\sigma_{\rm psf}$, above which the recovered fraction converges to the true value $f_{\rm tot}=0.5$.
Likelihood-based methods provide a powerful statistical framework for estimating the neutrino signal fraction $f_{\rm tot}$. However, it does not sufficiently address the impact of source confusion—a fundamental limitation in astronomical surveys that arises when the instrument’s angular resolution is too coarse to distinguish nearby sources (Barcons, 1992; Kurczynski and Gawiser, 2010; Jones et al., 2016). As also noted by Kowalski et al. (2025), stacking becomes unreliable when the average number of sources within a single PSF exceeds one, resulting in systematic underestimate of $f_{\rm tot}$.
@@ II.3 Mitigating Source Confusion in Analyses
Figure 2 summarizes the results for a fixed true value $f_{\rm tot}=0.5$. When the average separation satisfies $\bar{d}\geq 3\sigma_{\rm psf}$, the maximum-likelihood estimator reliably recovers the true $f$. In contrast, for denser source populations ($\bar{d}\leq 3\sigma_{\rm psf}$), confusion effects cause $f$ to be systematically underestimated. We note that the present analysis is based on the eFEDS AGN catalog, which covers a relatively limited sky area of $142\,\mathrm{deg}^{2}$. In regimes where the source density is low and the average separation becomes large, the finite survey area can amplify statistical fluctuations, resulting in increased scatter in the recovered $f_{\rm tot}$, as visible in the upper-left region of Figure 2. This effect is primarily driven by limited source statistics rather than a failure of the likelihood method itself. In future wide-area X-ray surveys with substantially larger sky coverage, this statistical limitation will be significantly alleviated, leading to more stable and accurate parameter recovery. This yields a practical condition for method applicability: the source population must be sufficiently sparse, such that the average spacing between sources exceeds approximately three times the PSF width. Importantly, we verify that this trend is valid for a wide range of values of $f_{\rm tot}$, not just for the specific case of $f_{\rm tot}=0.5$. In the following analyses, this empirical requirement, $\bar{d}\geq 3\sigma_{\rm psf}$, is adopted as a baseline condition to ensure that source confusion does not bias the likelihood inference.
Based on this empirical criterion, we set a minimum X-ray flux threshold $F_{\rm min}$ for all subsequent analyses, such that the selected AGN subset satisfies $\bar{d}\geq 3\sigma_{\rm psf}$. Since more luminous AGNs are more sparsely distributed on the sky, adjusting $F_{\rm min}$ provides a practical way to control the source density and to ensure that the selected population remains well resolved. This choice mitigates confusion-induced bias and preserves the validity of the maximum-likelihood inference. The resulting sample therefore represents only the relatively bright portion of the full AGN population. All reported values of $f_{\rm tot}$ in this work, including those shown in Figure 2, incorporate this correction, ensuring consistency with the full AGN population even though the likelihood analysis itself is restricted to the bright subset.
III Predictive Power for Future Neutrino Detectors
Next-generation neutrino observatories are expected to greatly enhance the capability to detect astrophysical neutrino sources. Proposed facilities such as IceCube-Gen2, HUNT, TRIDENT, NEON and Baikal-GVD, also promise enhanced exposure and angular resolution. These advances motivate us to examine how the improvements in angular resolution, event statistics, and energy threshold affect the prospects for diagnosing the neutrino–AGN correlation.
In this section, we consider a special scenario in which all astrophysical neutrinos originate from AGNs, corresponding to $\alpha=1$. This idealized case represents the maximal AGN contribution to the diffuse astrophysical neutrino flux and provides an upper limit on the expected signal in future detectors. Studying this scenario allows us to explore the full potential of next-generation instruments to constrain or detect neutrino–AGN associations under optimal conditions.